# Electromagnetic moments in the Sn-Gd region determined within the nuclear DFT

### v1

@@ 3 Results
In figures 12 and 13, which correspond to the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations in odd-$N$ and odd-$Z$ nuclei, respectively, we compared the calculated spectroscopic magnetic dipole moments with the experimental data. For the spin-aligned configuration, $\nu 11/2^{-}$, the agreement with the data is reasonably good [53]; however, the experimental trend is not reproduced in the Te isotopes. In contrast, the values and trends observed for the spin-anti-aligned configuration, $\pi 7/2^{+}$, do not align well with the data. It should be noted that in [11], the Sb experimental results were reproduced by adjusting the values of three effective $g$-factors in both shell-model and ab initio calculations. However, the physical rationale behind using these specific values remains unclear.
@@
As shown in figures 14 and 15, which compare the spectroscopic and intrinsic values of the calculated magnetic dipole moments, the restoration of rotational symmetry (AMP) is notable, with trends—particularly for $\nu 11/2$—and values altered.
@@ 4.1 Harmonic oscillator basis
The choice of the number of shells to include in the calculations depends on the convergence of results. Once a sufficient number of basis states are considered, increasing the basis size only slightly impacts the system’s properties. However, with a larger basis comes a more significant computational burden, making it essential to find the optimal set of states to balance precision and workload. To illustrate this analysis, in Figure 16, we plotted the energy of the oblate tag minimum of the ${11/2}^{-}$ state for the 167Gd nucleus as a function of different numbers of shells and HO frequencies. We can see that using a basis larger than $N_{0}=14$ has a minor effect on reducing the oblate minimum energy. For instance, in this example, the variation between $N_{0}=16$ and $N_{0}=18$ is 272 keV. Even more interestingly, at larger values of $N_{0}$, the results are nearly insensitive to the choice of $\hbar\omega_{0}$, cf. [56].
Figure 16: Ground state energies of the oblate tag minima in the ${11/2}^{-}$ state of 167Gd as functions of the HO basis frequency $\hbar\omega_{0}$.
Figure 17 presents the electric quadrupole moments (top panels) and magnetic dipole moments (bottom panels) determined at the oblate tag minimum (left panels) and prolate tag minimum (right panels) for different values of $N_{0}$ and $\hbar\omega_{0}$. We notice that the results obtained at $N_{0}\leq 14$ fluctuate significantly before stabilising. Specifically, at $N_{0}=16$, the standard deviations are in the range of $\Delta{Q}\simeq 0.03-0.05$ b and $\Delta{\mu}\simeq 0.002-0.003$$\mu_{N}$. To conclude, if we consider those deviations as typical uncertainties related to the basis truncation, the related error bars in all figures presented in this study would be invisible. For instance, the size of the symbols representing values of $Q$ ($\mu$) in figure 7 (figure 12) is 0.1 b (0.1 $\mu_{N}$), which is significantly larger than the numerical uncertainty estimated here.
Figure 17: (colour online) Spectroscopic electric quadrupole (upper panels) and magnetic dipole moments (lower panels) for the oblate (left panels) and prolate (right panels) minima of 167Gd using a spherical HO basis with different numbers of shells, $N_{0}$, and HO frequencies, $\omega_{0}$.
@@ 4.2 Pairing strength
Throughout this work, we observed a reasonable agreement with experimental data for the electric quadrupole moment. However, the calculated spectroscopic magnetic dipole moments of the ${\pi 7/2^{+}}$ configuration poorly reproduce the data. Since the magnetic dipole moments strongly depend on both the single-particle properties and the polarisation of the paired core, in this section, we critically evaluate the adjustment of the pairing strengths in the studied nuclei.
@@
Adjusting the pairing strengths would imply shifting the theoretical lines toward higher or lower values, which would not address the inconsistencies with the OES experimental values. Therefore, we conclude that the pairing strengths utilised in this work are reasonably calibrated, and there is no compelling reason to alter them. Although the OES analysis does not clarify the inconsistencies in our results for the magnetic dipole moments, it eliminates the pairing strengths as a potential factor. Additionally, this maintains the generality of DFT calculations without necessitating parameter adjustments for specific regions of the nuclear chart, which is one of the commendable features of this work.
@@ 4.3 Particle-number projection
In figure 19, we display the percentage deviations between the spectroscopic (AMP) magnetic dipole and electric quadrupole moments calculated with and without PNP. The PNP results were obtained using 9 Gauss-Tchebyshev integration nodes [37, 43] for both proton and neutron gauge angles. Tests were conducted for the $\nu 11/2^{-}$ configurations in odd-$N$ isotopes of gadolinium, yielding results that differ by only about 1% from those obtained without the PNP. This finding supports the sole use of AMP when determining electromagnetic moments. Given that the PNP increases the computational time by around two orders of magnitude (9$\times$9), this fact greatly enhances the feasibility of this work’s large-scale nuclear DFT calculations.
@@ 5 Conclusions and outlook
In this work, we have presented the nuclear DFT description of electric quadrupole and magnetic dipole moments in odd open-shell nuclei within the region of elements ranging from tin to gadolinium. Aligning the angular momenta along the axial symmetry axis, incorporating the signature and time-reversal breaking as well as time-odd mean fields, and restoring the rotational symmetry have proven to be critical elements of the description.
By tagging the quasiparticle states with the single-particle wave functions determined in 100Sn, we were able to trace two specific paired configurations of $\nu 11/2^{-}$ in odd-$N$ and $\pi 7/2^{+}$ in odd-$Z$ nuclei. We achieved reasonable agreement for both configurations between the calculated electric quadrupole moments and experimental data. Furthermore, for the parallel coupling of spin and orbital angular momenta in the $\nu 11/2^{-}$ configurations, the calculated magnetic dipole moments also align well with data, particularly in the Sn and Xe isotopes. In contrast, neither the values nor the trends in the experimental data could be reproduced for the antiparallel coupling in the $\pi 7/2^{+}$ configurations. We also examined the accuracy of the results against the truncation of the HO basis, strengths of the pairing interactions, and restoration of particle-number symmetry, concluding that none of these factors had a meaningful impact on the results. In comparison to restoring the particle-number symmetry and angular momentum symmetry for triaxial shapes, the methodology employed here reduced computational costs by four orders of magnitude and enabled efficient large-scale calculations.
A substantial disagreement exists between the calculated and experimental magnetic dipole moments of the $\nu 7/2^{+}$ configurations necessitating further investigation. On the one hand, attributing this disagreement to adjusted effective $g$-factors does not further scientific understanding. On the other hand, potential physical explanations for the disagreement identified here must be weighed against the agreement observed for the $\pi 11/2^{-}$ configurations. At this stage, we can argue that three promising avenues for further research exist. Firstly, one could aim to incorporate more sophisticated time-odd mean fields beyond those provided by the simple isovector spin-spin terms employed here. Secondly, in the region of the studied nuclei, the effects of triaxiality and collectivity may be significant. Thirdly, as suggested long ago (see, e.g., [59, 60]) and recently demonstrated in [61], the one-body magnetic dipole operator utilised in this study may require supplementation with terms generated by two-body meson-exchange currents. Work in these directions is the focus of our current research.
@@ Acknowledgments
This work was partially supported by STFC Grant Nos. ST/W005832/1, ST/P003885/1, ST/W50791X/1, and ST/V001035/1, and by a Leverhulme Trust Research Project Grant. We acknowledge the CSC-IT Center for Science Ltd., Finland, for allocating computational resources. This project was partly undertaken on the Viking Cluster, a high-performance compute facility provided by the University of York. We are grateful for computational support from the University of York High Performance Computing service, Viking, and the Research Computing team. We thank Grammarly® for its support with English writing.

### v2

@@ 3 Results
In figures 12 and 13, which correspond to the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations in odd-$N$ and odd-$Z$ nuclei, respectively, we compare the calculated spectroscopic magnetic dipole moments $\mu$ with experimental data. For the spin-aligned configuration, $\nu 11/2^{-}$, the agreement with the data is reasonably good [57]; however, the experimental trend is not reproduced in the Te isotopes. In contrast, the values and trends observed for the spin-anti-aligned configuration, $\pi 7/2^{+}$, do not align well with the data. It should be noted that in [7], the Sb experimental results were reproduced by adjusting the values of three effective $g$-factors in both shell-model and ab initio calculations. However, the physical rationale behind using these specific values remains unclear.
@@
As shown in figures 14 and 15, which compare the spectroscopic and intrinsic values of the calculated magnetic dipole moments, the restoration of rotational symmetry (AMP) is notable, with trends—particularly for $\nu 11/2$—and values altered accordingly.
@@ 4.1 Harmonic oscillator basis
The choice of the number of shells to include in the calculations depends on the convergence of results. Once a sufficient number of basis states are considered, increasing the basis size only slightly impacts the system’s properties. However, with a larger basis comes a more significant computational burden, making it essential to find the optimal set of states to balance precision and workload. To illustrate this analysis, in Figure 16, we plotted the energy of the oblate-tag minimum of the ${11/2}^{-}$ state for the 167Gd nucleus as a function of different numbers of shells and HO frequencies. We can see that using a basis larger than $N_{0}=14$ has a minor effect on reducing the oblate minimum energy. For instance, in this example, the variation between $N_{0}=16$ and $N_{0}=18$ is 272 keV. Even more interestingly, at larger values of $N_{0}$, the results are nearly insensitive to the choice of $\hbar\omega_{0}$, cf. [60].
Figure 16: Ground state energies of the oblate-tag minima in the ${11/2}^{-}$ state of 167Gd as functions of the HO basis frequency $\hbar\omega_{0}$.
Figure 17 presents the electric quadrupole moments $Q$ (top panels) and magnetic dipole moments $\mu$ (bottom panels) determined at the oblate-tag minimum (left panels) and prolate-tag minimum (right panels) for different values of $N_{0}$ and $\hbar\omega_{0}$. We notice that the results obtained at $N_{0}\leq 14$ fluctuate significantly before stabilising. Specifically, at $N_{0}=16$, the standard deviations are in the range of $\Delta{Q}\simeq 0.03-0.05$ b and $\Delta{\mu}\simeq 0.002-0.003$$\mu_{N}$. To conclude, if we consider those deviations as typical uncertainties related to the basis truncation, the related error bars in all figures presented in this study would be invisible. For instance, the size of the symbols representing values of $Q$ ($\mu$) in figure 7 (figure 12) is 0.1 b (0.1 $\mu_{N}$), which is significantly larger than the numerical uncertainty estimated here.
Figure 17: (colour online) Spectroscopic electric quadrupole moments $Q$ (upper panels) and magnetic dipole moments $\mu$ (lower panels) for the oblate (left panels) and prolate (right panels) minima of 167Gd using a spherical HO basis with different numbers of shells, $N_{0}$, and HO frequencies, $\omega_{0}$.
@@ 4.2 Pairing strength
Throughout this work, we observed a reasonable agreement with experimental data for the electric quadrupole moments $Q$. However, the calculated spectroscopic magnetic dipole moments $\mu$ of the ${\pi 7/2^{+}}$ configuration poorly reproduce the data. Since the magnetic dipole moments strongly depend on both the single-particle properties and the polarisation of the paired core, in this section, we critically evaluate the adjustment of the pairing strengths in the studied nuclei.
@@
Adjusting the pairing strengths would imply shifting the theoretical lines toward higher or lower values, which would not address the inconsistencies with the OES experimental values. Therefore, we conclude that the pairing strengths utilised in this work are reasonably calibrated, and there is no compelling reason to alter them. Although the OES analysis does not clarify the inconsistencies in our results for the magnetic dipole moments $\mu$, it eliminates the pairing strengths as a potential factor. Additionally, this maintains the generality of DFT calculations without necessitating parameter adjustments for specific regions of the nuclear chart, which is one of the commendable features of this work.
@@ 4.3 Particle-number projection
In figure 19, we display the percentage deviations between the spectroscopic (AMP) magnetic dipole moments $\mu$ and electric quadrupole moments $Q$ calculated with and without PNP. The PNP results were obtained using 9 Gauss-Tchebyshev integration nodes [41, 48] for both proton and neutron gauge angles. Tests were conducted for the $\nu 11/2^{-}$ configurations in odd-$N$ isotopes of gadolinium, yielding results that differ by only about 1% from those obtained without the PNP. This finding supports the sole use of AMP when determining electromagnetic moments. Given that the PNP increases the computational time by around two orders of magnitude (9$\times$9), this fact greatly enhances the feasibility of this work’s large-scale nuclear DFT calculations.
@@ 5 Conclusions and outlook
In this work, we presented the nuclear DFT description of electric quadrupole and magnetic dipole moments in odd open-shell nuclei, spanning the elements from tin to gadolinium. By aligning the angular momenta along the axial symmetry axis, incorporating the signature and time-reversal breaking as well as time-odd mean fields, and restoring rotational symmetry, we have proven these to be critical elements of the description.
By tagging the quasiparticle states with the single-particle wave functions determined in 100Sn, we were able to trace two specific paired configurations of $\nu 11/2^{-}$ in odd-$N$ and $\pi 7/2^{+}$ in odd-$Z$ nuclei. We achieved reasonable agreement for both configurations between the calculated electric quadrupole moments and experimental data. Furthermore, for the parallel coupling of spin and orbital angular momenta in the $\nu 11/2^{-}$ configurations, the calculated magnetic dipole moments also align well with data, particularly in the Sn and Xe isotopes. In contrast, neither the values nor the trends in the experimental data could be reproduced for the antiparallel coupling in the $\pi 7/2^{+}$ configurations. We also examined the precision of the results against the truncation of the HO basis, strengths of the pairing interactions, and restoration of particle-number symmetry, concluding that none of these factors had a meaningful impact on the results. In comparison to restoring the particle-number symmetry and angular momentum symmetry for triaxial shapes, the methodology employed here reduced computational costs by four orders of magnitude and enabled efficient large-scale calculations.
A substantial disagreement exists between the calculated and experimental magnetic dipole moments of the $\nu 7/2^{+}$ configurations, necessitating further investigation. On the one hand, attributing this disagreement to adjusted effective $g$-factors does not further scientific understanding. On the other hand, potential physical explanations for the disagreement identified here must be weighed against the agreement observed for the $\pi 11/2^{-}$ configurations. At this stage, we can argue that three promising avenues for further research exist. Firstly, one could aim to incorporate more sophisticated time-odd mean fields beyond those provided by the simple isovector spin-spin terms employed here. Secondly, in the region of the studied nuclei, the effects of triaxiality and collectivity may be significant. Thirdly, as suggested long ago (see, e.g., [63, 64]) and recently demonstrated in [65], the one-body magnetic dipole operator utilised in this study may require supplementation with terms generated by two-body meson-exchange currents. Work in these directions is the focus of our current research.
@@ Acknowledgments
This work was partially supported by STFC Grant Nos. ST/W005832/1, ST/P003885/1, ST/W50791X/1, and ST/V001035/1, and by a Leverhulme Trust Research Project Grant. We acknowledge the CSC-IT Center for Science Ltd., Finland, and the IFT Computer Center of the University of Warsaw, Poland, for allocating computational resources. This project was partly undertaken on the Viking Cluster, a high-performance compute facility provided by the University of York. We are grateful for computational support from the University of York High Performance Computing service, Viking, and the Research Computing team. We thank Grammarly® for its support with English writing.